OpenStudy (anonymous):

A soccer team played 160 games and won 65 percent of them. How many games did they win? how to solve?

OpenStudy (amistre64):

1% is a division by 100 160/100 = 1% they have 65 of those 1%s

OpenStudy (anonymous):

65 percent can be made into a fraction as 13/20. Now you can divide 160 by 20 which equals 8. Then multiply 8 by 13, which equals 104.

OpenStudy (anonymous):

How did you make 65 percent into a fraction?

OpenStudy (anonymous):

65% means 65/100. You can simplify 65/100 by finding the Greatest Common Factor (GCF) which is 5.

OpenStudy (amistre64):

well, since 1% is 1/100 and you have 65 of them ... 65(1/100) = 65/100 as a fraction

OpenStudy (anonymous):

Now you can divide 65 and 100 each by 5 and receive your new fraction of 13/20.

OpenStudy (anonymous):

oh! right! Thank you!